OpenStudy (anonymous):

How to simplify (n^2) /(n^2+2n+1) to be split into three terms?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first divide

OpenStudy (misty1212):

a nice gimmick is to write \[\frac{n^2}{n^2+2n+1}\] as \[\frac{n^2+2n+1-2n-1}{n^2+2n+1}=1-\frac{2n+1}{n^2+2n+1}\]

OpenStudy (misty1212):

partial fractions for the second part